a mixture of C2H2 and C3H8 is covered a certain volume at 80mmHg.mixture burn and make CO2 and H2O.if the pressure of CO2 is found 230mmHg at equal temperature and volume?
1
Expert's answer
2017-06-22T09:51:07-0400
C_2 H_2+(2+2/4) O_2→2CO_2+H_2 O
C_3 H_8+(3+8/4) O_2→3CO_2+4H_2 O
x+y=80
2x+3y=230
2(80-y)+3y=230
y=70.
The fraction of C3H8 in the mixture is
70/80 100%=87.5%.